OpenStudy (anonymous):

Suppose you invest $1050 at an annual rate of 5.5% compound continuously. How much will you have in the account after 5 years? round the solution to the nearest dollar

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

hey anyone know the answer?

OpenStudy (anonymous):

A=P(1+r/n)^tn P = principal amount (the initial amount you borrow or deposit - $1050) r = annual rate of interest (in decimal - .055) t = number of years the amount is deposited or borrowed for. n = number of times the interest is compounded per year

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

so 1050 (1.055) 5 ?

OpenStudy (anonymous):

and from that i got 1373

OpenStudy (anonymous):

P=I/RT

OpenStudy (anonymous):

You Got it :)

OpenStudy (anonymous):

1373? right

OpenStudy (anonymous):

Yes, Sir

OpenStudy (anonymous):

lol it was wrong